Given that a flux density of 1.2 is produced in a piece of steel by a magnetising force of 1250. Find the relative permeability of the steel under these conditions.
1
Expert's answer
2021-09-08T08:14:06-0400

"B=\mu \mu_0 H,"

"\mu=\frac B{\mu_0 H},"

"\mu=\frac {1.2}{4\pi\cdot 10^{-7}\cdot 1250}=764."